# From Mice to Trains: Amortized Bayesian Inference on Graph Data

### v3

@@ 1 Introduction
Amortized Bayesian Inference (ABI) (Gershman & Goodman, 2014; Papamakarios & Murray, 2018; Radev et al., 2023b; Ritchie et al., 2016) provides a compelling solution to a series of problems like repeated posterior inference in simulator-based models, real-time updating in sequential decision-making, and scalable analyses across large collections of related datasets: By jointly training multiple neural networks to approximate posterior distributions, ABI enables rapid and repeated inference even when the likelihood is intractable. ABI underpins several strands of simulation-based inference, including neural likelihood estimation and neural posterior estimation (Papamakarios et al., 2019; Mittal et al., 2025; Radev et al., 2023a). Despite notable successes in various fields like uncertainty aware surrogate modeling (Scheurer et al., 2025), Bayesian mixture models (Pakman & Paninski, 2018; Kucharský & Bürkner, 2025), Bayesian multilevel models (Habermann et al., 2025; Arruda et al., 2025; 2024; Wu et al., 2024) and experimental design (Huang et al., 2024; Kennamer et al., 2023), existing ABI pipelines are not tailored to graphs and face new challenges when working with graph-structured data.
@@ 2.1 Problem setting
In our Bayesian formulation, we model parameters $\theta$ and observed data $D$ jointly via $p(\theta,D)=p(\theta)\,p(D\mid\theta)$. Here, $D$ denotes an observed graph, and the latent parameters $\theta$ govern properties the structural and statistical properties of this graph.
@@ 2.3.5 Graph Transformer
2.4 Metrics
In the experiments (Section 3), we instantiate the ABI pipeline with each of the proposed architectures as the summary network across a range of problem settings. To enable a fair comparison and to quantify posterior quality, we report three complementary metrics — calibration, posterior contraction, and recovery — assessing, respectively, coverage reliability, concentration of uncertainty, and correspondence between true parameters and posterior summaries.
@@ 3 Experiments
We conduct three experiments to compare neural architectures used as summary networks. The first experiment is a controlled toy problem, in which we evaluate 12 summary-network variants on amortized posterior estimation. The second experiment is a biologically motivated case study that models the interaction network of free-ranging mice. The third experiment considers a logistics application and evaluates model performance in a neural likelihood estimation setting. All code and material can be found on GitHub [^1] https://github.com/sjedhoff/ABI-graph-paper.git.
@@ 3.1.2 Training architecture
We use the proposed ABI framework to enable posterior inference on the parameters $\theta$. The summary network takes the simulated graph with the adjacency matrix and the feature vector $(\tilde{A},X)$ as input and produces a fixed-length summary vector. In this case study, we use a 16-dimensional summary, which exceeds the parameter dimension of four and follows the heuristic that the summary dimension should not be smaller than the number of target parameters (Li et al., 2025). We evaluate four summary-network architectures: a Graph Convolutional Network (GCN), Deep Sets, a Graph Transformer, and a Set Transformer. For each architecture, we test three pooling mechanisms: simple mean pooling as a baseline, the permutation-invariant layer from Deep Sets (see Eq. 8), and, as the most expressive option, pooling via multi-head attention (PMA) (see Eq. 18). The inference network takes the summary $s=h(D)$ as input and outputs an approximate posterior $p_{\phi}(\theta\mid s)$ over $\theta$. We use a coupling flow with spline transformations build with 6 invertible layers. Each model is trained online for 250 epochs, with 100 batches per epoch and a batch size of 32.
@@ 3.1.3 Results
Figure 2 summarizes the experimental results. For better visualization, the three parameters $\pi_{AA},\pi_{BB}$, and $\pi_{AB}$ are presented together in the first row labeled $\pi$. Table 2 in Appendix A gives additional insights about the results, such as the value of the training loss at the final-epoch and the number of trainable parameters for each summary architecture.
@@
Figure 3 compares a representative run of the Graph Convolutional Network with mean pooling and the Set Transformer with multi-head attention pooling in terms of recovery and calibration. The recovery plot (left) shows estimates for $\pi_{AB}$ as a representative example of the three $\pi$ parameters. For both $\pi_{AB}$ and $\lambda$, the Graph Convolutional Network exhibits poorer recovery than the Set Transformer. Vertical lines denote 95% credible intervals, which are noticeably narrower for the Set Transformer, indicating sharper estimates. The Graph Convolutional Network appears to learn mainly the prior distribution, a common failure mode, whereas the Set Transformer learns a meaningful posterior, which can be seen by a correlation of 0.873 and 0.847 for $\pi_{AB}$ and $\lambda$ respectively of posterior-median and true parameter value. The ECDF calibration plots (right panel of Figure 3) provide a diagnostic of posterior calibration based on simulation-based ranks, as described in Section 2.4. Both the Graph Convolutional Network and the Set Transformer calibrate the parameter adequately, as their rank ECDFs lie within the 95% confidence bands.
@@ 3.2.1 Simulator setup
In this case study, we simulate a group of 30 mice. Each mouse is initialized with 5 taxa drawn from a pool of 20 possible taxa. Because taxa are mouse-specific, we encode their abundances in the node-feature matrix $X\in\mathbb{R}^{30\times 20}$, where each row corresponds to a mouse and each column to a taxon. Active taxa are represented by zero; present taxa are stored as relative amount that sum to $100\%$ per mouse. Taxa whose amount falls below a presence threshold, here 0.0001%, are removed from a mouse’s microbiome. A taxon is also removed if it is not exchanged for four days in a row.
@@ 3.2.2 Training architecture
As summary networks, we evaluate the four architectures introduced in Section 2.3: Deep Sets (with invariant pooling), a Graph Convolutional Network (with invariant pooling), a Graph Transformer, and a Set Transformer (both using multi-head attention pooling). We fix the summary dimension to 16. The inference network is a coupling flow with spline-based transformations build with 6 layers. For each observation horizon (5, 10, and 30 days), we generate 50 00050\,000 simulations for training and an additional $1000$ simulations for validation. All models use dropout with rate $0.05$ and apply early stopping based on the validation loss to mitigate overfitting.
@@ 3.2.3 Results
The results across summary-network architectures and the three observation horizons are summarized in Figure 5. Table 3 in Appendix B shows the exact median values of the metrics together with the final validation loss. The network-density parameter $\delta$ is recovered reliably: in most settings, recovery, measured as the correlation between posterior medians and ground-truth values, exceeds 0.9. Posterior contraction likewise indicates that the inferred posteriors for $\delta$ are suitably concentrated. However, the calibration score $\ell_{\gamma}$ reveals that, on average, none of the models achieves well-calibrated posteriors for $\delta$.
@@ 4 Summary
Across three empirical case studies, the paper evaluates graph-aware ABI using several permutation-invariant summary networks. In the first case study, a controlled toy simulator generates undirected graphs with type-specific baseline edge probabilities and an additional triadic-closure parameter. Most architectures infer the baseline probabilities accurately, yielding high parameter recovery and strong posterior contraction. In contrast, the triadic-closure parameter reflected higher-order, local structure, making it substantially harder to identify; among the tested models, the Deep Sets, Graph Transformer and Set Transformer achieve high parameter recovery and posterior contraction. Achieving good calibration remains challenging for architectures, which achieve good parameter recovery. Second, in a biological simulator of weighted social interaction networks among mice driving microbiome exchange, the global network density is inferred reliably but was not well calibrated, while the exchange factor was more difficult. Across all considered summary-network architectures, parameter recovery declines as the observation horizon increased, consistent with reduced identifiability once the microbiome dynamics approach a steady state. The Set Transformer again outperforms the other three architectures in terms of recovery and posterior contraction. Interestingly, the best calibration for the network-density parameter is achieved by Deep Sets and the Graph Transformer, and for the exchange-factor parameter by the Graph Convolutional Network, even though the Graph Convolutional Network performs worst overall in recovery and contraction. Third, in a train-scheduling simulator defined on a fixed rail-section graph, the task is to perform neural likelihood estimation for the trains’ total travel times. Here, the Set Transformer combined with attention pooling produced well-calibrated posteriors over the full travel-time distributions: posterior medians closely track the ground truth, and the inferred densities capture the characteristic right-skewness and frequent multimodality induced by stochastic delays and traffic conflicts.
@@ Appendix A Toy Example: Additional Results
Appendix B Mice interaction network: Additional Results

### v4

@@ 1 Introduction
Amortized Bayesian Inference (ABI) (Gershman & Goodman, 2014; Papamakarios & Murray, 2018; Radev et al., 2023b; Ritchie et al., 2016; Gloeckler et al., 2024) provides a compelling solution to a series of problems like repeated posterior inference in simulator-based models, real-time updating in sequential decision-making, and scalable analyses across large collections of related datasets: By jointly training multiple neural networks to approximate posterior distributions, ABI enables rapid and repeated inference even when the likelihood is intractable. ABI underpins several strands of simulation-based inference, including neural likelihood estimation and neural posterior estimation (Papamakarios et al., 2019; Mittal et al., 2025; Radev et al., 2023a). Despite notable successes in various fields like uncertainty aware surrogate modeling (Scheurer et al., 2025), Bayesian mixture models (Pakman & Paninski, 2018; Kucharský & Bürkner, 2025), Bayesian multilevel models (Habermann et al., 2025; Arruda et al., 2025; 2024; Wu et al., 2024) and experimental design (Huang et al., 2024; Kennamer et al., 2023), existing ABI pipelines are not tailored to graphs and face new challenges when working with graph-structured data. The limited prior work in this direction either focuses on spatial data (Sainsbury-Dale et al., 2025) or addresses narrow, problem-specific settings (Stillman et al., 2023), leaving a general framework for graph-structured ABI largely undeveloped.
@@ 2.1 Problem setting
In our Bayesian formulation, we model parameters $\theta$ and observed data $D$ jointly via $p(\theta,D)=p(\theta)\,p(D\mid\theta)$. Here, $D$ denotes an observed graph, and the latent parameters $\theta$ govern the structural and statistical properties of this graph.
@@ 2.3.5 Graph Transformer
2.4 Evaluation metrics
In the experiments (Section 3), we instantiate the ABI pipeline with each of the proposed architectures as the summary network across a range of problem settings. To enable a fair comparison and to quantify posterior quality, we report three complementary metrics — calibration, posterior contraction, and recovery — assessing, respectively, coverage reliability, concentration of uncertainty, and correspondence between true parameters and posterior summaries (Lueckmann et al., 2021; Hermans et al., 2022). In all experiments, metrics are computed on unseen simulated test data.
@@ 3 Experiments
We conduct three experiments to compare neural architectures used as summary networks. The first experiment is a controlled toy problem, in which we evaluate 12 summary-network variants on amortized posterior estimation. The second experiment is a biologically motivated case study that models the interaction network of free-ranging mice. The third experiment considers a logistics application and evaluates model performance in a neural likelihood estimation setting. Since none of the three experiments have a tractable likelihood, direct comparison with non-SBI baselines such as MCMC is not possible. To give an indication of how our ABI approach compares to such a baseline, we present a simplified example with a tractable likelihood in Appendix C. All code and material for the experiments can be found on GitHub [^1] https://github.com/sjedhoff/ABI-graph-paper.git.
@@ 3.1.2 Training architecture
We use the proposed ABI framework to enable posterior inference on the parameters $\theta$. The summary network takes the simulated graph with the adjacency matrix and the feature vector $(\tilde{A},X)$ as input and produces a fixed-length summary vector. In this case study, we use a 16-dimensional summary, which exceeds the parameter dimension of four and follows the heuristic that the summary dimension should not be smaller than the number of target parameters (Li et al., 2025). We evaluate four summary-network architectures: a Graph Convolutional Network (GCN), Deep Sets, a Graph Transformer, and a Set Transformer. For each architecture, we test three pooling mechanisms: simple mean pooling as a baseline, the permutation-invariant layer from Deep Sets (see Eq. 8), and, as the most expressive option, pooling via multi-head attention (PMA) (see Eq. 18). The inference network takes the summary $s=h(D)$ as input and outputs an approximate posterior $p_{\phi}(\theta\mid s)$ over $\theta$. We use a coupling flow with spline transformations built with 6 invertible layers. Each model is trained online for 250 epochs, with 100 batches per epoch and a batch size of 32.
@@ 3.1.3 Results
Figure 2 summarizes the experimental results. For better visualization, the three parameters $\pi_{AA},\pi_{BB}$, and $\pi_{AB}$ are presented together in the first row labeled $\pi$. Table 2 in Appendix A provides additional details on the results, including the final training loss and the number of trainable parameters for each summary network architecture. Table 3 in the same appendix reports training times across different dataset sizes, and Figure 11 illustrates how dataset size affects model performance.
@@
Figure 3 compares a representative run of the Graph Convolutional Network with mean pooling and the Set Transformer with multi-head attention pooling in terms of recovery and calibration. The recovery plot (left) shows estimates for $\pi_{AB}$ as a representative example of the three $\pi$ parameters. For both $\pi_{AB}$ and $\lambda$, the Graph Convolutional Network exhibits poorer recovery than the Set Transformer. Vertical lines denote 95% credible intervals, which are noticeably narrower for the Set Transformer, indicating sharper estimates. The Graph Convolutional Network appears to learn mainly the prior distribution, a common failure mode, whereas the Set Transformer learns a meaningful posterior, which can be seen by a correlation of 0.873 and 0.847 for $\pi_{AB}$ and $\lambda$ respectively of posterior-median and true parameter value. The ECDF calibration plots (right panel of Figure 3) provide a diagnostic of posterior calibration based on simulation-based ranks, as described in Section 2.4. Both the Graph Convolutional Network and the Set Transformer calibrate the parameter adequately, as their rank ECDFs lie within the 95% confidence bands. Model performance can be assessed not only through parameter-based metrics, but also through data-dependent quantities. Appendix B presents data-dependent SBC results for this experiment.
@@ 3.2.1 Simulator setup
In this case study, we simulate a group of 30 mice. Each mouse is initialized with 5 taxa drawn from a pool of 20 possible taxa. Because taxa are mouse-specific, we encode their abundances in the node-feature matrix $X\in\mathbb{R}^{30\times 20}$, where each row corresponds to a mouse and each column to a taxon. Active taxa are represented by zero; present taxa are stored as relative amounts that sum to $100\%$ per mouse. Taxa whose amount falls below a presence threshold, here 0.0001%, are removed from a mouse’s microbiome. A taxon is also removed if it is not exchanged for four days in a row.
@@ 3.2.2 Training architecture
As summary networks, we evaluate the four architectures introduced in Section 2.3: Deep Sets (with invariant pooling), a Graph Convolutional Network (with invariant pooling), a Graph Transformer, and a Set Transformer (both using multi-head attention pooling). We fix the summary dimension to 16. The inference network is a coupling flow with spline-based transformations built with 6 layers. For each observation horizon (5, 10, and 30 days), we generate 50 00050\,000 simulations for training and an additional $1000$ simulations for validation. All models use dropout with rate $0.05$ and apply early stopping based on the validation loss to mitigate overfitting.
@@ 3.2.3 Results
The results across summary-network architectures and the three observation horizons are summarized in Figure 5. Table 4 in Appendix D shows the exact median values of the metrics together with the final validation loss. The network-density parameter $\delta$ is recovered reliably: in most settings, recovery, measured as the correlation between posterior medians and ground-truth values, exceeds 0.9. Posterior contraction likewise indicates that the inferred posteriors for $\delta$ are suitably concentrated. However, the calibration score $\ell_{\gamma}$ reveals that, on average, none of the models achieves well-calibrated posteriors for $\delta$.
@@
3.3 Real-world data
To demonstrate the method’s applicability to real-world data, we construct a graph using data from Raulo et al. (2023). Social associations between mice were quantified using the adjusted simple ratio index, which measures the frequency with which two mice were observed at the same location within a 12-hour window (see Raulo et al. (2024) for details). Although this spatiotemporal co-occurrence is not a direct measure of social interaction, it is a widely used proxy for social association among individuals (Albery et al., 2021; Sah et al., 2019).
From the microbiome data, we focused on anaerobic non-sporeformers, as these taxa are the primary drivers of microbiome transmission through social networks (Raulo et al., 2024). To keep the analysis computationally feasible, we retained only the taxa with the highest variance in occurrence across mice, excluding taxa that are nearly omnipresent or nearly absent, as these carry little discriminatory information. This resulted in a network of 177 mice and 52 taxa.
Since the graph size differs from those used in previous training, we retrain the model on simulated data that more closely resembles the true dataset. The new training data is generated with adapted hyperparameters and with the observed social association network held fixed. Because the network is fixed, its density is no longer a free parameter, leaving the exchange factor as the sole parameter of interest. For the ABU model, we use a Set Transformer with PMA aggregation as the summary network, motivated by its strong performance in earlier experiments. Since inference is performed on a single parameter, we adopt flow matching (Liu et al., 2022) as the inference network. Both networks are trained together offline on 10 00010\,000 simulated datasets.
Figure 7 shows the recovery and calibration evaluated on a simulated test dataset, alongside the posterior density of the exchange factor inferred from the real graph. Recovery on the test data is very high. The calibration shows a tendency toward overestimation, though this may partly be an artifact of the high recovery, as small deviations become detectable when the overall fit is already very good. The posterior density (right panel) is strongly concentrated compared to the uniform prior on (0.05,0.95), with a posterior mean of 0.144 and a posterior standard deviation of 0.009.
Figure 7: Recovery (left) and calibration (center) of the exchange factor $\alpha$ evaluated on simulated test data for the real-world mice model. The right panel shows the posterior distribution of $\alpha$ inferred from the real-world dataset, displayed as a histogram with a kernel density estimate overlay.
We further evaluate the model using posterior predictive checks. For each posterior draw, we simulate microbiome data using the simulator and summarize the output via the mean and standard deviation of the Jaccard index. The Jaccard index captures the proportion of microbial taxa shared between a pair of mice relative to the total detected across both, providing an intuitive measure of transmission signal (Raulo et al., 2024; 2021).
Figure 8 plots the mean Jaccard index against its standard deviation. The blue points represent prior predictive samples, obtained by simulating data across 100 equally spaced draws from the prior, and thus illustrate the range of summary statistic combinations the simulator can produce. The red points correspond to posterior draws conditioned on the real-world dataset, shown in green. While the mean and standard deviation of the real-world dataset each fall within the prior predictive distribution individually, their combination lies outside the region the simulator can reproduce. Comparing the real-world value to the posterior draws, the standard deviation is well matched, but the mean Jaccard index of the real-world data is substantially higher than that of the posterior simulations.
These results indicate that the model captures the variance in the data reasonably well, but fails to recover the correct location. This discrepancy does not stem solely from the trained model; it primarily reflects a misspecified simulator. The posterior predictive checks reveal a clear simulation gap between the data the simulator can generate and the structure present in the real-world dataset. As this paper focuses on the application of ABI to graph-structured data rather than on capturing the full biological complexity of mouse gut microbiome transmission, we do not attempt to close this gap here.
Figure 8: Posterior predictive check for the real-world mouse microbiome dataset. Mean Jaccard index (x-axis) versus standard deviation of the Jaccard index (y-axis) for simulated and observed data. Blue points represent prior predictive samples, illustrating the range of summary statistic combinations attainable by the simulator. Red points correspond to posterior predictive samples conditioned on the real-world dataset. The green point marks the observed value for the real-world dataset.
@@ 4 Summary
Across three empirical case studies, the paper evaluates graph-aware ABI using several permutation-invariant summary networks. In the first case study, a controlled toy simulator generates undirected graphs with type-specific baseline edge probabilities and an additional triadic-closure parameter. Most architectures infer the baseline probabilities accurately, yielding high parameter recovery and strong posterior contraction. In contrast, the triadic-closure parameter reflected higher-order, local structure, making it substantially harder to identify; among the tested models, the Deep Sets, Graph Transformer and Set Transformer achieve high parameter recovery and posterior contraction. Achieving good calibration remains challenging for architectures, which achieve good parameter recovery. Second, in a biological simulator of weighted social interaction networks among mice driving microbiome exchange, the global network density is inferred reliably but was not well calibrated, while the exchange factor was more difficult. Across all considered summary-network architectures, parameter recovery declines as the observation horizon increases, consistent with reduced identifiability once the microbiome dynamics approach a steady state. The Set Transformer again outperforms the other three architectures in terms of recovery and posterior contraction. Interestingly, the best calibration for the network-density parameter is achieved by Deep Sets and the Graph Transformer, and for the exchange-factor parameter by the Graph Convolutional Network, even though the Graph Convolutional Network performs worst overall in recovery and contraction. Third, in a train-scheduling simulator defined on a fixed rail-section graph, the task is to perform neural likelihood estimation for the trains’ total travel times. Here, the Set Transformer combined with attention pooling produced well-calibrated posteriors over the full travel-time distributions: posterior medians closely track the ground truth, and the inferred densities capture the characteristic right-skewness and frequent multimodality induced by stochastic delays and traffic conflicts.
@@ Appendix A Toy Example: Additional Results
Table 3: Training time in minutes for different training dataset sizes across the four summary network architectures in the toy example from Section 3.1. All networks use the invariant aggregation layer (Deep Sets and GCN) or PMA aggregation (Graph Transformer and Set Transformer). Once trained, all architectures generate posterior samples near-instantaneously, requiring approximately 0.1 seconds for 1000 draws. All reported runtimes were measured on an Intel Core Ultra 9 185H CPU (2.30 GHz) with 32 GB of RAM.
Training Data Size DeepSets GCN GraphTransformer SetTransformer
3,200 4.94 3.38 9.04 13.78
6,400 10.59 6.41 29.47 50.67
12,800 36.40 11.90 58.40 73.95
32,000 127.88 41.10 175.56 294.58
Figure 11: Recovery of the parameters $\pi=\{\pi_{AA},\pi_{BB},\pi_{AB}\}$ and $\lambda$ for four summary network architectures as a function of training dataset size. All networks use either the invariant aggregation layer (Deep Sets and GCN) or PMA aggregation (Graph Transformer and Set Transformer). For all architectures except the GCN, recovery improves only marginally with increasing dataset size, whereas the GCN shows substantial gains as more training data becomes available.
Appendix B Toy Example: Data-dependent SBC
In the main text, we discussed the performance of the networks measured with SBC on model parameters, but calibration can also be evaluated at the level of the data itself. This approach can be understood as a form of posterior predictive checking; since we operate in a simulation-based setting, we follow Modrák et al. (2025) and employ data-dependent test quantities for SBC. A natural and well-established choice is the model likelihood, but in many settings where SBI is applied, the likelihood is intractable and therefore unavailable as a test quantity. Since our data are graphs, we require test quantities that summarize graph structure. We opt for four such measures: the spectral gap, edge density, degree assortativity, and the global clustering coefficient.
The spectral gap is the difference between the first and second eigenvalue of the normalized graph Laplacian. A larger spectral gap indicates a well-connected graph that is difficult to disconnect, whereas a small spectral gap is characteristic of graphs with pronounced community structure or near-disconnection (Fiedler, 1973).
Edge density is the fraction of possible edges that are present in the graph. It provides a basic summary of graph sparsity and directly corresponds to the connection probability in Erdős–Rényi random graphs (Newman, 2010).
Degree assortativity is the Pearson correlation coefficient between the degrees of adjacent node pairs. Positive values indicate that high-degree nodes tend to connect to other high-degree nodes, a pattern commonly observed in social networks, while negative values indicate that hubs tend to connect to low-degree nodes, as is typical in biological and technological networks (Newman, 2002).
The global clustering coefficient, also referred to as transitivity, measures the fraction of connected node triplets that close into triangles: $3\times\text{triangles}/\text{triplets}$. It captures the tendency of a network to form tightly knit local groups (Watts & Strogatz, 1998).
To obtain values of the introduced metrics for a fitted model, the following steps are performed: A graph sampled from the simulator serves as the test graph, and posterior draws of the parameters are obtained using the fitted model. Each posterior draw is then passed into the simulator to generate a new graph, for which the metrics described above are computed and compared to those calculated from the original test graph.
Figure 12 shows the results of the data-dependent SBC for the toy example from Section 3.1, covering all twelve summary networks and the four data-dependent test metrics. The left column displays the recovery, i.e., the correlation between the true metric value and the median of the metrics computed from the posterior draws. For the Deep Sets, Graph Transformer, and Set Transformer architectures, recovery is very high across all four data-dependent metrics. In particular, the Deep Sets and Set Transformer achieve a recovery of 1 for both edge density and global clustering coefficient. The GCN struggles to attain good recovery when using mean or invariant aggregation, whereas the GCN with PMA aggregation performs comparably to the other three network types. Calibration of these metrics is challenging, particularly when recovery is very high, as even minor miscalibration becomes detectable. For edge density, the Deep Sets and Set Transformer models achieve near-perfect recovery but fail to calibrate well — the values of $\ell_{\gamma}$ for the Set Transformer collapse to -$\infty$. For degree assortativity, all networks achieve good calibration, with $\ell_{\gamma}>0$ throughout.
Overall, most networks show satisfactory SBC performance with the chosen data-dependent test quantities. The calibration difficulties can be attributed to the extremely high recovery of the metrics and should therefore not be interpreted as evidence of poor calibration.
Figure 12: Results for the toy example across four summary-network architectures with each three different aggregation layers. We report recovery (higher is better) and calibration ($\ell_{\gamma}$; values > 0 indicate good calibration) for the metrics spectral gap, edge density, degree assortativity and the global clustering coefficient. Markers (circles/triangles/squares) indicate the median across five runs, and error bars span the minimum to maximum values. Markes at the very bottom of the plot represent the value of -$\infty$.
Appendix C Toy example: Comparison to MCMC baseline
To compare our proposed approach against an MCMC baseline, a tractable likelihood is required. However, in all three experiments — including the toy example — this is not the case. To nonetheless provide an indication of how the ABI approach compares to MCMC, we introduce a simplified version of the toy example in which only the parameters $\pi_{AA}$, $\pi_{BB}$, and $\pi_{AB}$ are used to simulate a graph. The triadic closure parameter $\lambda$ is removed to yield a tractable likelihood.
We train a model on this reduced simulator using a Set Transformer with PMA aggregation as the summary network and a coupling flow with spline transformation as the inference network within BayesFlow (Kühmichel et al., 2026). As a baseline, we fit the same model via MCMC, specifically HMC (Neal, 2011) implemented in Stan (Stan Development Team, 2024). Online training of the BayesFlow model for 150 epochs with 100 batches of size 32 per epoch (480,000 datasets in total) took approximately 38.85 minutes. Drawing 500 posterior samples for each of 500 test datasets was then nearly instantaneous, requiring only 0.92 seconds in total. For the MCMC baseline, four separate chains are run for each of the 500 test graphs, which took 33.43 minutes in total.
Figure 13 shows calibration ECDF plots for both the BayesFlow model and the MCMC baseline. Both approaches achieve good calibration for $\pi_{AA}$, $\pi_{BB}$, and $\pi_{AB}$. The log-likelihood, however, is only well calibrated under MCMC; the BayesFlow model struggles to reach comparable calibration for this quantity. This discrepancy is expected, as the BayesFlow networks are trained using a simulation-based loss rather than the likelihood-based objective that underpins MCMC.
Figure 13: Calibration ECDF plots for the BayesFlow model using a Set Transformer (with PMA aggregation) as the summary network compared to a MCMC baseline in the simplified version of the toy experiment. In addition to the parameters of the simulator $\pi_{AA},\pi_{BB}$ and $\pi_{AB}$, the calibration of the log-likelihood is shown.
Figure 14 shows the recovery of the three parameters, with MCMC posterior medians on the x-axis and the corresponding BayesFlow posterior medians on the y-axis. Error bars represent 95% credible intervals. The posterior medians from BayesFlow closely match those from MCMC, and the widths of the error bars are similar in both directions, indicating that the two approaches yield comparable posterior uncertainty.
Taken together, these results demonstrate that the BayesFlow model closely approximates the MCMC baseline. It should be noted, however, that MCMC is only applicable when an analytical likelihood is available, which was not the case in the experiments presented in Section 3. Furthermore, while the total compute times are comparable — 38 minutes of training for BayesFlow versus 33 minutes for MCMC across 500 test graphs — the BayesFlow model has a decisive practical advantage: once trained, it produces posterior samples for any new graph in under a second, whereas MCMC must be run from scratch for each new observation.
Figure 14: Recovery of the parameters $\pi_{AA},\pi_{BB}$ and $\pi_{AB}$ for the MCMC baseline vs the model fitted with BayesFlow (summary network: Set Transformer with PMA aggregation; inference network: coupling flow with spline transformation). The horizontal and vertical lines correspond to the 95 % credible intervals.
Appendix D Mice interaction network: Additional Results